40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 5 19:38:52 2012, jamie, jenne, ayaka, den, Update, Alignment, more alignment woes 
    Reply  Tue Nov 6 16:38:37 2012, jenne, jamie, ayaka, manasa, Update, Alignment, Alignment back under control again 
       Reply  Tue Nov 6 17:07:04 2012, jamie, Update, Alignment, AS and REFL tmp6oTENk.pngtmplamEtZ.png
       Reply  Tue Nov 6 17:22:51 2012, Manasa, Jamie, Update, Alignment, Alignment- POY and oplevs 
          Reply  Wed Nov 7 00:10:38 2012, Jenne , Update, Alignment, Alignment- POY and oplevs. photos. 
             Reply  Wed Nov 7 07:11:10 2012, rana, Update, Alignment, Alignment- POY and oplevs. photos. 
             Reply  Wed Nov 7 09:09:02 2012, Steve, Update, Alignment, Alignment- 9amNov7w.png
                Reply  Wed Nov 7 15:17:15 2012, Steve, Update, Alignment, steering option with pico motor? IMG_1795.JPG
Message ID: 7675     Entry time: Tue Nov 6 17:22:51 2012     In reply to: 7673     Reply to this: 7677
Author: Manasa, Jamie 
Type: Update 
Category: Alignment 
Subject: Alignment- POY and oplevs 

Right now, Manasa, Jamie and Ayaka are doing some finishing touches work, checking that POY isn't clipping on OM2, the second steering mirror after the SRM, and they'll confirm that POX comes out of the chamber nicely, and that POP is also still coming out (by putting the green laser pointer back on that table, and making sure the green beam is co-aligned with the beam from PR2-PR3.  Also on the list is checking the vertex oplevs.  Steve and Manasa did some stuff with the ETM oplevs yesterday, but haven't had a chance to write about it yet.

We were trying to check POY alignment using the green laser in the reverse direction (outside vacuum to in-vac) . The green laser was installed along with a steering mirror to steer it into the ITMY chamber pointing at POY.

We found that the green laser did follow the path back into the chamber perfectly; it was clipping at the edge of POY. To align it to the center of POY (get a narrower angle of incidence at the ITMY), the green laser had to be steered in at a wider angle of incidence from the table. This is now being limited by the oplev steering optics on the table. We were not able to figure out the oplev path on the table perfectly; but we think we can find a way to move the oplev steering mirrors that are now restricting the POY alignment.

The oplev optics will be moved once we confirm with Jenne or Steve.

 

[Steve, Manasa]

We aligned the ETM oplevs yesterday. We confirmed that the oplev beam hit the ETMs. We checked for centering of the beam coming back at the oplev PDs and the QPDsums matched the values they followed before the vent.

Sadly, they have to be checked once again tomorrow because the alignment was messed up all over again yesterday.

ELOG V3.1.3-